A doctor orders 20 grams of a 52% solution of a certain medicine. The pharmacist has only bottles of 40% and bottles of 70% solution. How much of each must he use to obtain the 20 grams of the 52% solution?
Answer by scott8148(6628)

let x="grams of 40%", so 20-x="grams of 70%" ... x(40%)+(20-x)(70%)=20(52%)

.4x+1.4-.7x=10.4